OpenStudy (anonymous):

Find the quotient: -72 ÷ (-12) A. -7 B. -6 C. 6 D. 8

OpenStudy (anonymous):

6

OpenStudy (anonymous):

-72 /-12=6 negative will be positive

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (anonymous):

if u r satisfied plz give medal

OpenStudy (anonymous):

suree (:

OpenStudy (unklerhaukus):

\[\frac{-72 }{-12}=\frac{\cancel {-6}\times12}{\cancel {-6}\times2}=\frac{\cancel2\times6}{\cancel2}=6 \]